Operations Report – November 30th

SOL: 04

Name of person filing report: Munkh-Erdene "Muggi" Altankhuyag

Non-nominal systems: No non-operational system

Notes on non-nominal systems: Generator issue has been fixed and all systems back to plugged including supplemental light in the Hab.


Spirit rover used: not-used

Hours: (before EVA)

Beginning charge: (Before EVA)

Ending charge: (On return from EVA, before recharging)

Currently charging: unplugged

Opportunity rover used:


Beginning charge:

Ending charge:

Currently charging: unplugged

Curiosity rover used: the rover is used for EVA-04, and we

Hours: 259.3

Beginning charge: 100%

Ending charge: 33%

Currently charging: unplugged

Perseverance rover used: The rover is used for EVA-04

Hours: 278.4

Beginning charge: 100%

Ending charge: 29%

Currently charging: unplugged

General notes on rovers: The rovers have been taken to Brahe Hwy1572, Perseverance rover’s parking brake wasn’t operational.

Summary of Hab operations: added 110 gallons of water

WATER USE: 65.27 gallons

Water (static tank): 288.729 gallons

Static tank pipe heater (on or off): on

Static tank heater (On or off) on

Toilet tank emptied: yes

Summary of internet: Internet in the hab was disabled until 6:00 pm.

Summary of suits and radios: all radios are operational, and suits are operational but not charging, will be pulled in this night

Summary of GreenHab operations: Systems are operational.

WATER USE: 0.79 gallons

Heater: On

Supplemental light: will be operated from 10pm to 2 am

Harvest: 0.09lb

Summary of ScienceDome operations: The Dome systems are operational, EVA’s Cap Comm used the dome.

Dual split: Heat or AC, Off

Summary of RAM operations: RAM wasn’t used.

Summary of any observatory issues: nothing to report

Summary of health and safety issues: nothing to report

Questions, concerns and requests to Mission Support: nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site